OpenStudy (anonymous):

A CD tower is shaped like a cylinder and holds CDs with a diameter of 4.75 inches. If the CD tower is 12 inches tall, what is its volume in cubic inches? Use 3.14 for pi. Round your answer to the nearest hundredth. 212.54 cubic inches 231.56 cubic inches 850.16 cubic inches 867.34 cubic inches can anyone help explain this to me ? the formula is V=πr^2h

OpenStudy (anonymous):

what is the radius

OpenStudy (anonymous):

remember, the radius is half the diameter

OpenStudy (anonymous):

Well you know the height= 12in You don't know the radius but you can find it since the diameter is given radius=1/2 diameter diameter=4.75 r=1/2(4.75in) r=? h=12in V=πr^2h

OpenStudy (anonymous):

212.54 the formula is the area of your base times your height. So it would be 4.75 divided by 2 in order to get your radius, the you would square it and multiply it by 3.14 and then multiply it by 12:)

OpenStudy (anonymous):

Btw @deeezzzz the volume of the cylinder is 3.14 x r( squared) x height. You dont multiply the height by 2